OpenStudy (anonymous):

Which shows the numbers in order from least to greatest? 1.38,1 3/7, 1.4

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

1 3/7=10/7=1.42 1.38, 1.42,1.4

OpenStudy (midhun.madhu1987):

1.38 , 1.428 , 1.4 So, which is the greatest among these?

OpenStudy (gorv):

can u arrange it

OpenStudy (anonymous):

so its 1.38 1.4 142

OpenStudy (anonymous):

right

OpenStudy (gorv):

yeah right .well done :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

wait its 1.37 1.38 1.4

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

u was rigth earlier

OpenStudy (anonymous):

but there is o anser like that

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

thn u made a mistake in typin the Q

OpenStudy (anonymous):

can you help me with one more? @gorv

OpenStudy (gorv):

plzz post new one